# On Bayes factor functions

### v2

@@ 3 Comparisons of alternative classes of prior distributions
    A normal-moment prior (Johnson, Pramanik and Shudde, 2023),
    $j(\lambda,{\,|\,}\tau^{2})=\frac{\lambda^{2}}{\sqrt{2\pi}\tau^{3}}\exp\left(-\frac{\lambda^{2}}{2\tau^{2}}\right).$
When the $t$ test arises from a test of whether a normal mean is zero based on a random sample $\{X_{i}\}\ \mbox{i.i.d.}\ N(\mu,\sigma^{2})$, $i=1,\dots,n$, then $\lambda=\sqrt{n}\mu/\sigma$. The Cauchy prior (2) then corresponds to the JZS prior, specified as
@@
where $\chi^{-2}_{1}$ denotes an inverse chi-squared distribution on one degree of freedom (Rouder et al., 2009, 2012). Alternative (3) corresponds to a g-prior, also with a non-informative prior imposed on $\sigma^{2}$ (Liang et al., 2008). Alternative (4) is a special case of the class of normal moment priors described in Datta et al. (2025).
EPP plots for these alternative prior densities appear in Fig. 3. EPP values are indexed by the IQR of each alternative prior density. For example, the IQR of the Cauchy density is $2r$, while the IQR of the normal prior is $1.35\sqrt{g}$. The values of $\tau$ that correspond to a specified IQR of the inverse-moment and normal-moment priors were determined numerically. The values in Fig. 3 reflect the EPP associated with each alternative prior as its scale parameters is determined so that it assigns 0.5 probability to the same interval around the null value.
Figure 3: EPP plots for $t$-test statistics with 15 degrees-of-freedom. The horizontal axis provides the IQR that each prior density assigns around the null value of the non-centrality parameter (i.e., $\lambda=0$).
Fig. 3 highlights several contrasts between the EPP associated with the four alternative prior densities. When the IQR of the priors exceeds about 5.0, the non-local prior densities provide significantly higher EPP than the local priors. Recalling that the noncentrality of the $t$ density is $\sqrt{n}\mu/\sigma$, an IQR of 5.0 corresponds to, for example, a sample size of 100 and a standardized effect of 0.5 (i.e., a medium effect size (Cohen, 2013)).
For small values of the prior IQR, the Cauchy and inverse-moment priors provide slightly better EPP, but, on average, all priors provide unconvincing evidence for the true model. Perhaps surprisingly, the JZS prior and the inverse-moment priors provide approximately the same EPP for small prior IQRs.
@@
Figure 5: EPP plots of normal moment priors of different orders, $r$, for testing whether the noncentrality parameter of a $t_{15}$ statistic is 0. For comparison, the EPP plot for inverse moment alternative priors with shape parameter $\nu=9$ is provided by the black curve.
Figure 6: EPP plots for testing the non-centrality parameter $\lambda$ of a $\chi^{2}_{5}$ random variable. Alternative prior densities on $\lambda$ include an inverse gamma with shape 4.5, a gamma prior with shape 4.5, a squared-Cauchy density, and an exponential prior.
EPP plots can also be used to compare non-local priors. Fig. 5 displays an EPP plot for normal moment priors (Datta et al., 2025), along with an inverse-moment prior with shape parameter $\nu=9$.
Fig.6 displays an EPP plot for testing whether the non-centrality parameter $\lambda$ of a $\chi^{2}_{5}$ statistic is 0. Four alternative prior densities are included in the comparison. These include an exponential prior, a gamma prior with shape parameter 4.5, an inverse gamma prior with shape parameter $\nu/2=4.5$, and a squared-Cauchy prior. The latter is motivated by a Cauchy prior on a normal mean, which implies a squared-Cauchy prior on a chi-squared non-centrality parameter. The density of a squared Cauchy is
$f(x)=\frac{1}{\pi r\sqrt{x}(1+x/r^{2})}.$ (13)
The squared-Cauchy prior provides the strongest support for the true model when concentrated near 0, but comparatively weaker support for more disperse choices of its scale parameter. Except for small differences near 0, the inverse gamma prior demonstrates higher EPP than the other priors that have the same median.
@@ 4.1 Selection of the shape parameter, $\nu$
EPP plots analogous to Figure 3 can also be constructed for $z$-tests, $t$-tests with arbitrary degrees of freedom, and test statistics based on $\chi^{2}$ and $F$ distributions. In principle, such plots can be combined with subjective prior information to guide the selection of a class of alternative prior distributions to construct BFFs. Alternatively, information from replicated study designs or meta-analyses may inform this selection; this possibility is further explored in Section 6.
Both inverse-moment and inverse-gamma prior families are characterized by shape and scale parameters. Focusing first on inverse-moment priors, EPP plots can inform the choice of the shape parameter $\nu$. Figure 7 illustrates such a strategy. As in previous figures, the horizontal axis represents the interquartile range (IQR) of the alternative prior, which is determined by the scale parameter $\tau$.
This plot reveals similar features to those in Figure 3. For small IQR values, choosing $\nu=1$ yields the highest EPP. However, as the IQR increases, the prior mass near the origin decreases more rapidly for higher values of $\nu$, as illustrated in Figure 1.
In the absence of subjective prior information or data from replication studies, these findings suggest that setting $\nu=1$ is a conservative default. This choice generally provides higher EPP than larger values of $\nu$ when the alternative prior is concentrated near the null, while also outperforming commonly used local priors as the noncentrality parameter increases.
A potential drawback of selecting $\nu=1$ is that its improved EPP near the null hypothesis comes at the cost of reduced EPP for alternative priors that assign greater mass to larger effect sizes. Moreover, the gains in EPP near the null tend to correspond to EPP values in the range of 0.3–0.7, which may be considered inconclusive from a scientific perspective.
In many areas of the social and biological sciences, prior knowledge concerning typical standardized effect sizes is available. Investigators often focus on detecting effects within the standardized range 0.2 (small) to 0.8 (large), with 0.5 denoting a medium effect size (Cohen, 2013). Consequently, we recommend selecting $\nu$ so that the alternative prior on the noncentrality parameter $\lambda$ places a specified proportion of its mass—say, $\gamma\approx 0.9$—within the range corresponding to standardized effects in $(0.2,0.8)$, assuming the prior mode for $\omega$ is 0.5.
@@
Figure 7: Plots of EPP for various inverse-moment priors as a function of their IQR.
@@ 4.2 Indexing Bayes factor functions by scale parameters
The noncentrality parameter of a test statistic provides one possibility, the IQR another. When non-local alternative priors are used, indexing the BFF by a standardized effect size is also possible. For inverse-moment and inverse gamma priors, this can be accomplished by setting the scale parameter $\tau$ so that the mode of the prior density coincides with a specified standardized effect (Johnson, Pramanik and Shudde, 2023).
@@
$r(\omega_{1})=\underset{\lambda}{\arg\max}\ i(\lambda{\,|\,}\tau_{\omega_{1},\nu}).$ (14)
@@
Based on this procedure, Table 2 provides formulae for setting the scale parameter $\tau$ so that the modes of inverse-moment and inverse gamma densities are placed at values of the noncentrality parameter that correspond to a hypothesized standardized effect $\omega_{1}$. Given the shape parameter $\nu$, inverse-moment densities $i(\lambda{\,|\,}\tau,\nu)$ are imposed on $z$ and $t$ tests, and inverse gamma densities $ig(\lambda{\,|\,}\tau,\nu/2)$ are assumed for $\chi^{2}$ and $F$ tests. This table is similar to Table 1 of (Johnson, Pramanik and Shudde, 2023), which describe $\tau^{2}$ values for $r^{th}$ normal-moment and gamma prior densities.
@@
Table 2: For normal mean tests, $\bar{x}$ is the sample mean of $n$ independent $N(\mu,\sigma^{2})$ deviates, while $\bar{x_{i}}$ and $n_{i}$, $i=1,2$, are the corresponding quantities in two sample tests. The variance is $\sigma^{2}$. Standard deviations are denoted by $s$ and are the pooled estimate in the two-sample $t$ test. For multinomial/Poisson tests, $f(\bm{\theta})$ maps an $s\times 1$ vector $\bm{\theta}$ into a $k\times 1$ probability vector, $k$ the number of cells. The degrees-of-freedom $\delta=k-s-1$; $p_{i}$ and $n_{i}$ are cell probabilities and counts, and $n$ is the sum of cell counts. In the Linear Model, the alternative hypothesis is $\mathbf{A}\bm{\beta}=\mathbf{a}$, where $\mathbf{A}$ is a $k\times p$ matrix of rank $k$, $\bm{\beta}$ is a $p$ vector of regression coefficients, and $\mathbf{a}$ is a $k$ vector. The quantities $RSS_{0}$ and $RSS_{1}$ denote the residual sum-of-squares under the null and alternative hypotheses. For the Likelihood Ratio test, $l(\cdot)$ denotes the likelihood function for ${\bm{\theta}}=({\bm{\theta}}_{r},{\bm{\theta}}_{s})$. The $k\times 1$ subvector ${\bm{\theta}}_{r}$ equals ${\bm{\theta}}_{r0}$ under the null hypothesis. The MLE of $\bm{\theta}$} under $H_{1}$ is $\hat{{\bm{\theta}}}$, and the MLE of ${\bm{\theta}}_{s}$ under $H_{0}$ is $\hat{{\bm{\theta}}}_{s}$. The matrix ${\bf L}^{-1}$ represents the Cholesky decomposition of the covariance matrix $\bf C$} for the tested parameters, scaled to a single observation. The final column provides the value of $\tau_{\omega_{1},\nu}$ that places the mode of the prior density at the non-centrality parameter specified in column 5.
@@ 5.1 Two-sample t statistics
The two-sample $t$-test is among the most commonly used test statistics in the social sciences. To illustrate BFFs based upon it, suppose $x_{ij}$}, $j=1,2$, $i=1,\dots,n_{j}$ represent conditionally independent draws from $N(\mu_{j},\sigma^{2})$ distributions, and, following Table 2, we define the standardized effect as $\omega=(\mu_{1}-\mu_{2})/\sigma$, $s^{2}$ as the pooled estimate of the variance, and $\bar{x}_{j}$ as the sample mean of each group. The $t$ statistic is then defined as
@@
For $p\in(0.5,0.025,0.005,0.0025)$ let $q_{1-p,\delta}$ denote the $(1-p)^{\rm th}$ quantiles of standard $t$ statistics on $\delta=n_{1}+n2-2$ degrees-of-freedom, i.e.,
@@
Under these assumptions, BFFs for quantiles of a two-sample $t$ statistic with $n_{1}=n_{2}=15$ and an inverse moment prior with shape parameter $\nu=9$ are displayed Fig. 8. This figure provides a concise summary of evidence in favor of a range of alternative hypotheses based on $t$ statistics of varying degrees of statistical significance. For instance, the maximum Bayes factor in favor of the alternative hypothesis for any combination of $q_{1-p,28}$ and $\omega_{1}$ was 24.1, corresponding to $q_{.9975,28}=3.05$ and an inverse-moment prior centered on a standardized effect $\omega_{1}=0.68$ and noncentrality parameter $\lambda_{1}=15\omega_{1}/\sqrt{30}=1.86$. At the other extreme, decisive evidence in favor of the null hypothesis was obtained for $q_{0.5,28}=t_{28}=0$ for all alternative priors centered on effect sizes greater than $\omega_{1}=0.83$.
Figure 8: BFFs for $t_{28}$ test statistics definedd with inverse-moment priors. The horizontal bands indicate the strength of evidence provided by the Bayes factors for the given value of $\omega_{1}$. $\rm P_{1}$ denotes positive evidence ($\rm BF_{10}\in(3-10)$), $\rm S_{1}$ strong evidence (10-30), $\rm VS_{1}$ very strong (30-100), or $\rm D_{1}$ decisive evidence. The reciprocal values provide evidence in favor of the null hypothesis. For example, $\rm P_{0}$ denotes positive evidence for the null hypothesis, or $\rm BF_{10}\in(1/10-1/3)$.
Figure 9: BFFs for $t_{28}$ test statistics deefined with JZS priors. The labeling of horizontal bands is similar to Fig. 8, as is the scaling of the vertical axis.
For comparison, Fig. 9 provides a similar plot for JSZ priors (Rouder et al., 2009), indexed by the Cauchy scale parameter $r$. For two-sample $t$ statistics, the noncentrality parameter is $\sqrt{n_{1}n_{2}/(n_{1}+n_{2})}\omega$, implying the prior on the $t$ non-centrality parameter is a Cauchy centered on 0 with scale parameter $r\sqrt{n_{1}n_{2}/(n_{1}+n_{2})}$. The BFFs in this plot do not indicate strong evidence in favor of the alternative hypothesis for any values of $r$ for any of the $t_{28}$ quantiles. Nor do they provide strong evidence in favor of the null hypothesis for commonly recommended values of $r$ (i.e., ($\sqrt{2}/2,2,\sqrt{2}$)).

### v3

@@ 3 Comparisons of alternative classes of prior distributions
    Normal-moment priors (Johnson et al.,, 2023), with $r=1$ in equation (3).
When a $t$ test arises from a test of whether a normal mean is zero based on a random sample $\{X_{i}\}\ \mbox{i.i.d.}\ N(\mu,\sigma^{2})$, $i=1,\dots,n$, then $\lambda=\sqrt{n}\mu/\sigma$. The Cauchy prior (2) then corresponds to the Jeffreys–Zellner–Siow (JZS) prior, specified as
@@
where $\chi^{-2}_{1}$ denotes an inverse chi-squared distribution on one degree of freedom (Rouder et al.,, 2009, 2012). Alternative (3) corresponds to a g-prior, also with a non-informative prior imposed on $\sigma^{2}$ (Liang et al.,, 2008). Alternative (4) is a special case of the class of normal moment priors described in Datta et al., (2025).
$\rm D_{EP}$} plots for these alternative prior densities versus the null hypothesis appear in Fig. 3. $\rm D_{EP}$} values are indexed by the IQR of each alternative prior density. For example, the IQR of the Cauchy density is $2r$, while the IQR of the normal prior is $1.35\sqrt{g}$. The values of $\tau$ that correspond to a specified IQR of the inverse-moment and normal-moment priors were determined numerically. The values in Fig. 3 reflect the $\rm D_{EP}$} associated with each alternative prior versus the null as its scale parameter is determined so that it assigns 0.5 probability to the same interval around the null value.
Figure 3: $\rm D_{EP}$} plots for $t$-test statistics with 15 degrees-of-freedom. The horizontal axis provides the IQR that each prior density assigns around the null value of the non-centrality parameter (i.e., $\lambda=0$).
Fig. 3 highlights several contrasts between the $\rm D_{EP}$} associated with the four alternative prior densities. When the IQR of the priors exceeds about 5.0, the non-local prior densities provide significantly higher $\rm D_{EP}$} than the local priors. Recalling that the noncentrality of the $t$ density is $\lambda=\sqrt{n}\mu/\sigma$, an IQR of 5.0 corresponds to, for example, a sample size of 100 and a standardized effect of 0.5 (i.e., a medium effect size (Cohen,, 2013)).
For small values of the prior IQR, the Cauchy and inverse-moment priors provide slightly better $\rm D_{EP}$}, but, on average, all priors provide unconvincing evidence for the true model. Perhaps surprisingly, the JZS prior and the inverse-moment priors provide approximately the same $\rm D_{EP}$} for small prior IQRs.
@@
Figure 5: $\rm D_{EP}$} plots of normal moment priors of different orders, $r$, for testing whether the noncentrality parameter of a $t_{15}$ statistic is 0. For comparison, the $\rm D_{EP}$} plot for inverse moment alternative priors with shape parameter $\nu=9$ is provided by the black curve.
Figure 6: $\rm D_{EP}$} plots for testing the non-centrality parameter $\lambda$ of a $\chi^{2}_{5}$ random variable. Alternative prior densities on $\lambda$ include an inverse gamma with shape 4.5, a gamma prior with shape 4.5, a squared-Cauchy density, and an exponential prior.
$\rm D_{EP}$} plots can also be used to compare non-local priors versus null hypotheses. Fig. 5 displays an $\rm D_{EP}$} plot for normal moment priors (Datta et al.,, 2025), along with an inverse-moment prior with shape parameter $\nu=9$. Fig. 6 displays an $\rm D_{EP}$} plot for testing whether the non-centrality parameter $\lambda$ of a $\chi^{2}_{5}$ statistic is 0. Four alternative prior densities are included in the comparison. These include an exponential prior, a gamma prior with shape parameter 4.5, an inverse gamma prior with shape parameter $\nu/2=4.5$, and a squared-Cauchy prior. The latter is motivated by a Cauchy prior on a normal mean, which implies a squared-Cauchy prior on a chi-squared non-centrality parameter (Rouder et al.,, 2012). The density of a squared Cauchy is
$f(x)=\frac{1}{\pi r\sqrt{x}(1+x/r^{2})}.$
The squared-Cauchy prior provides the strongest support for the true model when concentrated near 0, but comparatively weaker support for more disperse choices of its scale parameter. Except for small differences near 0, the inverse gamma prior demonstrates higher $\rm D_{EP}$} than the other priors that have the same median.
@@ 4.1 Selection of the shape parameter, $\nu$
$\rm D_{EP}$} plots analogous to Figure 3 can also be constructed for $z$-tests, $t$-tests with arbitrary degrees of freedom, and test statistics based on $\chi^{2}$ and $F$ distributions. In principle, such plots can be combined with subjective prior information to guide the selection of a class of alternative prior distributions to construct BFFs. Alternatively, information from replicated study designs or meta-analyses may inform this selection; this possibility is further explored in Section 6.
Both inverse-moment and inverse-gamma prior families are characterized by shape and scale parameters. Focusing first on inverse-moment priors, $\rm D_{EP}$} plots can inform the choice of the shape parameter $\nu$. Figure 7 illustrates such a strategy. As in previous figures, the horizontal axis represents the interquartile range (IQR) of the alternative prior, which is determined by the scale parameter $\tau$.
This plot reveals similar features to those in Figure 3. For small IQR values, choosing $\nu=1$ yields the highest $\rm D_{EP}$}. However, as the IQR increases, the prior mass near the origin decreases more rapidly for higher values of $\nu$, as illustrated in Figure 1.
In the absence of subjective prior information or data from replication studies, these findings suggest that setting $\nu=1$ is a conservative default. This choice generally provides higher $\rm D_{EP}$} than larger values of $\nu$ when the alternative prior is concentrated near the null, while also outperforming commonly used local priors as the noncentrality parameter increases.
A potential drawback of selecting $\nu=1$ is that its improved $\rm D_{EP}$} near the null hypothesis comes at the cost of reduced $\rm D_{EP}$} for alternative priors that assign greater mass to larger effect sizes. Moreover, the gains in $\rm D_{EP}$} near the null tend to correspond to $\rm D_{EP}$} values in the range of 0.3–0.7, which may be considered inconclusive from a scientific perspective.
In many areas of the social and biological sciences, prior knowledge concerning typical standardized effect sizes is available. Investigators often focus on detecting effects within the standardized range 0.2 (small) to 0.8 (large), with 0.5 denoting a medium effect size (Cohen,, 2013). Consequently, we recommend selecting $\nu$ so that the alternative prior on the noncentrality parameter $\lambda$ places a specified proportion of its mass—say, $\gamma\approx 0.9$—within the range corresponding to standardized effects in $(0.2,0.8)$, assuming the prior mode for $\omega$ is 0.5.
@@
Figure 7: Plots of $\rm D_{EP}$} for various inverse-moment priors as a function of their IQR.
@@ 4.2 Indexing Bayes factor functions by scale parameters
The noncentrality parameter of a test statistic provides one potential index, the IQR another. When non-local alternative priors are used, indexing the BFF by a standardized effect size is also possible. For inverse-moment and inverse gamma priors, this can be accomplished by setting the scale parameter $\tau$ so that the mode of the prior density coincides with a specified standardized effect (Johnson et al.,, 2023).
@@
$r(\omega_{1})=\underset{\lambda}{\arg\max}\ i(\lambda{\,|\,}\tau_{\omega_{1},\nu}).$
@@
Based on this procedure, Table 1 provides formulae for setting the scale parameter $\tau$ so that the modes of inverse-moment and inverse gamma densities are placed at values of the noncentrality parameter that correspond to a hypothesized standardized effect $\omega_{1}$. Given the shape parameter $\nu$, inverse-moment densities $i(\lambda{\,|\,}\tau,\nu)$ are imposed on $z$ and $t$ tests, and inverse gamma densities $ig(\lambda{\,|\,}\tau,\nu/2)$ are assumed for $\chi^{2}$ and $F$ tests. This table is similar to Table 1 of (Johnson et al.,, 2023; Datta et al.,, 2025), which describe $\tau^{2}$ values for normal-moment and gamma prior densities.
@@
For normal mean tests, $\bar{x}$ is the sample mean of $n$ independent $N(\mu,\sigma^{2})$ deviates, while $\bar{x}_{j}$ and $n_{j}$, $j=1,2$, are the corresponding quantities in two sample tests. The variance is $\sigma^{2}$. Standard deviations are denoted by $s$ and are the pooled estimate in the two-sample $t$ test. For multinomial/Poisson tests, $f(\bm{\theta})$ maps an $s\times 1$ vector $\bm{\theta}$ into a $k\times 1$ probability vector, $k$ the number of cells. The degrees-of-freedom $\delta=k-s-1$; $p_{i}$ and $n_{i}$ are cell probabilities and counts, and $n$ is the sum of cell counts. In the Linear Model, the alternative hypothesis is $\mathbf{A}\bm{\beta}=\mathbf{a}$, where $\mathbf{A}$ is a $k\times p$ matrix of rank $k$, $\bm{\beta}$ is a $p$ vector of regression coefficients, and $\mathbf{a}$ is a $k$ vector. The quantities $RSS_{0}$ and $RSS_{1}$ denote the residual sum-of-squares under the null and alternative hypotheses. For the Likelihood Ratio test, $l(\cdot)$ denotes the likelihood function for ${\bm{\theta}}=({\bm{\theta}}_{r},{\bm{\theta}}_{s})$. The $k\times 1$ subvector ${\bm{\theta}}_{r}$ equals ${\bm{\theta}}_{r0}$ under the null hypothesis. The MLE of $\bm{\theta}$} under $H_{1}$ is $\hat{{\bm{\theta}}}$, and the MLE of ${\bm{\theta}}_{s}$ under $H_{0}$ is $\hat{{\bm{\theta}}}_{s}$. The matrix ${\bf L}^{-1}$ represents the Cholesky decomposition of the covariance matrix $\bf C$} for the tested parameters, scaled to a single observation. The final column provides the value of $\tau_{\omega_{1},\nu}$ that places the mode of the prior density at the non-centrality parameter specified in column 5.
@@ 5.1 Two-sample t statistics
The two-sample $t$-test is among the most commonly used test statistics in the social sciences (Cohen,, 1994; Maxwell and Delaney,, 2004). To illustrate BFFs based upon it, suppose $x_{ij}$}, $j=1,2$, $i=1,\dots,n_{j}$ represent conditionally independent draws from $N(\mu_{j},\sigma^{2})$ distributions, and, following Table 1, we define the standardized effect as $\omega=(\mu_{1}-\mu_{2})/\sigma$, $s^{2}$ as the pooled estimate of the variance, and $\bar{x}_{j}$ as the sample mean of each group. The $t$ statistic is then defined as
@@
For $p\in(0.5,0.025,0.005,0.0025)$ let $q_{1-p,\delta}$ denote the $(1-p)^{\rm th}$ quantiles of standard $t$ statistics on $\delta=n_{1}+n_{2}-2$ degrees-of-freedom, i.e.,
@@
Under these assumptions, BFFs for quantiles of a two-sample $t$ statistic with $n_{1}=n_{2}=15$ and an inverse moment prior with shape parameter $\nu=9$ are displayed Fig. 8. This figure provides a concise summary of evidence in favor of a range of alternative hypotheses based on the observation of quantiles of $t$ statistics drawn from the null distribution. For instance, $t_{15}$ statistics equal to 2.05, 2.76, and 3.05 (two-sided p-values of 0.05, 0.01, 0.005) all provide positive evidence for the alternative hypothesis for certain $\omega_{1}$ values, with the latter statistics also providing strong evidence for $\omega_{1}$ values around 0.6. The maximum Bayes factor in favor of the alternative hypothesis was 24.1, achieved for $q_{1-p,28}=3.05$ with an inverse-moment prior centered on a standardized effect $\omega_{1}=0.68$, corresponding to a noncentrality parameter of $\lambda_{1}=15\omega_{1}/\sqrt{30}=1.86$. At the other extreme, decisive evidence in favor of the null hypothesis was obtained for $q_{0.5,28}=t_{28}=0$ for all alternative priors centered on effect sizes greater than $\omega_{1}=0.83$.
Figure 8: BFFs for $t_{28}$ test statistics defined with inverse-moment priors. The horizontal bands indicate the strength of evidence provided by the Bayes factors for the given value of $\omega_{1}$. $\rm P_{1}$ denotes positive evidence ($\rm BF_{10}\in(3-10)$) and $\rm S_{1}$ strong evidence (1030)(10-30). The reciprocal values provide evidence in favor of the null hypothesis. $\rm VS_{0}$ denotes very strong evidence for the null hypothesis ($\rm BF_{10}\in(1/30-1/100)$), and $\rm D_{0}$ decisive evidence.
Figure 9: BFFs for $t_{28}$ test statistics defined with JZS priors. The labeling of horizontal bands is similar to Fig. 8, as is the scaling of the vertical axis.
For comparison, Fig. 9 provides a similar plot for JSZ priors (Rouder et al.,, 2009), indexed by the Cauchy scale parameter $r$. For two-sample $t$ statistics, the noncentrality parameter is $\sqrt{n_{1}n_{2}/(n_{1}+n_{2})}\omega$, implying the prior on the $t$ non-centrality parameter is a Cauchy centered on 0 with scale parameter $r\sqrt{n_{1}n_{2}/(n_{1}+n_{2})}$. The BFFs in this plot do not indicate strong evidence in favor of the alternative hypothesis for any values of $r$ for any of the $t_{28}$ quantiles. Nor do they provide strong evidence in favor of the null hypothesis for commonly recommended values of $r$ (i.e., ($\sqrt{2}/2,2,\sqrt{2}$)).